# Project page: https://zhang-tao-whu.github.io/projects/DVIS_DAQ

### v4

@@ Abstract
Modern video segmentation methods adopt object queries to perform inter-frame association and demonstrate satisfactory performance in tracking continuously appearing objects despite large-scale motion and transient occlusion. However, they all underperform on newly emerging and disappearing objects that are common in the real world because they attempt to model object emergence and disappearance through feature transitions between background and foreground queries that have significant feature gaps. We introduce Dynamic Anchor Queries (DAQ) to shorten the transition gap between the anchor and target queries by dynamically generating anchor queries based on the features of potential candidates. Furthermore, we introduce a query-level object Emergence and Disappearance Simulation (EDS) strategy, which unleashes DAQ’s potential without any additional cost. Finally, we combine our proposed DAQ and EDS with DVIS [71] to obtain DVIS-DAQ. Extensive experiments demonstrate that DVIS-DAQ achieves a new state-of-the-art (SOTA) performance on five mainstream video segmentation benchmarks. Code and models are available at https://github.com/SkyworkAI/DAQ-VS.
@@ 1 Introduction
However, these query-based methods exhibit a notable limitation: they tend to underperform when dealing with newly emerging or disappearing objects, even with recent state-of-the-art (SOTA) methods [71, 21]. As shown in the left part of Fig. 1, these methods only achieve less than 45% recall ratio on emerging and disappearing objects. As illustrated in the middle part of Fig. 1, they often fail to detect newly emerging objects. And they may incorrectly track a different, similar object when the original object disappears. Given that many objects appear and disappear over time in real-world long videos, it is imperative to improve these methods to handle objects’ emergence and disappearance more effectively, thereby meeting the requirements of real-world application requirements.
@@ 2 Related Work
Video Instance Segmentation (VIS). Earlier works [66, 35, 67] extend image instance segmentation methods by adding tracking heads and learning the feature association. With the rise of vision transformer [8, 54, 16], current SOTA video instance methods [54, 22, 10, 59, 23, 60, 68, 33, 37, 28, 18] adopt query-based designs. In particular, Video K-Net [34] and IDOL [60] directly learn the query association embeddings via contrastive learning. After that, association-based methods [66, 23, 60, 68, 33, 37], accomplish the tracking of video targets between adjacent frames or adjacent video clips through object matching. Meanwhile, several models [22, 10, 59] directly predict 3D volumes of video instances across times in a semi-online manner. On the other hand, propagation-based methods [21, 28, 18, 71] delegate the tracking problem to the transformer decoder. They take the object queries outputted from the previous frame or video clip as input and iteratively optimize them to segment the objects in the current frame. Note that these methods can also be applied for video panoptic segmentation (VPS) and video semantic segmentation (VSS) tasks [48, 40, 39, 55, 24, 78] by adding queries for segmenting the stuff. However, both poorly handle challenges such as new-emerging objects and target disappearance. Our approach improves the model’s ability to address these challenges by employing dynamic anchor queries to manage the emergence and disappearance of objects.
@@
Universal Video Segmentation. Current efforts in the field of video segmentation directly draw upon the design of universal image segmentation models [56, 73, 10, 12, 64, 13, 53] to develop universal video segmentation models [34, 26, 33, 1, 71, 31, 30, 58], achieving comparable or even superior performance to specialized video segmentation methods. Video K-Net [34] and TubeFormer [26] are the first to unify all three video segmentation tasks. TarVIS [1] extends the concept of mask classification from image segmentation to video segmentation, accomplishing universal video segmentation via one shared model. DVIS [71] receives frame object query outputs from Mask2Former [10] for universal video segmentation. Recently, several works [32, 62] unified both image and video segmentation in one model. However, these methods are still unaware of object emergence and disappearance problems, which are significant for long videos.
@@
Video Object Tracking. Object tracking [70, 50, 43, 74, 51, 63, 42, 17, 14, 7, 3, 75] is a crucial task in VIS and VPS [76, 24, 25, 4, 29, 19], and many works adopt the tracking-by-detection paradigm [6, 27, 61, 77, 46, 57, 52, 38]. These methods divide the task into two sub-tasks, where an object detector first detects objects and then associates them using a tracking algorithm. Recent works [44, 15, 49, 2] also perform clip-wise segmentation and tracking. However, the former [44, 15] only focuses on single-object mask tracking, while the latter [49] considers global tracking via clip-level matching. Our work belongs to mask-based tracking, which is orthogonal to these existing tracking methods.
@@ 4.3 Ablation Studies and Analysis
Qualitative results comparison. In Fig. 7, we present the visual comparison with our baseline. In this challenging scene, there are new-emerging, disappearing, and reappearing video objects. Our method accurately addresses all these scenarios, whereas DVIS suffers from issues including missed new-emerging objects and ID switches. Although DVIS detected object 1 in frame 2, it incorrectly switched the ID of another object to that of object 1. Regarding the reappearance of objects 1 and 2 in Frame 6, our method successfully maintained continuous tracking. In Frame 7, DVIS incorrectly switched the IDs of two nearby disappearing objects to objects 5 and 6, whereas our method correctly handled these disappearing cases.
@@ 5 Conclusion
Limitations. We have significantly enhanced the capability of current query-based methods to handle newly emerging and disappearing objects. However, we introduced an additional tracker specifically to address object disappearance. While the additional computational cost is minimal, it undoubtedly introduces more hyperparameters. In the future, we will explore using a single model to effectively handle continuously appearing, newly emerging, and disappearing objects. Additionally, our model reaches its performance peak after 160k training iterations. Although this issue is prevalent in other methods, such as GenVIS and DVIS++, we believe there is ample room for optimization to enhance the model’s training efficiency.
@@ 0.A.4 More Qualitative Results
Fig. A1 presents qualitative results in complex scenes. To fully demonstrate the performance of our method in challenging scenes such as fast motion and occlusions, we also provide supplementary videos for reference.

### v5

@@ Abstract
Modern video segmentation methods adopt object queries to perform inter-frame association and demonstrate satisfactory performance in tracking continuously appearing objects despite large-scale motion and transient occlusion. However, they all underperform on newly emerging and disappearing objects that are common in the real world because they attempt to model object emergence and disappearance through feature transitions between background and foreground queries that have significant feature gaps. We introduce Dynamic Anchor Queries (DAQ) to shorten the transition gap between the anchor and target queries by dynamically generating anchor queries based on the features of potential candidates. Furthermore, we introduce a query-level object Emergence and Disappearance Simulation (EDS) strategy, which unleashes DAQ’s potential without any additional cost. Finally, we combine our proposed DAQ and EDS with the previous method, DVIS, to obtain DVIS-DAQ. Extensive experiments demonstrate that DVIS-DAQ achieves a new state-of-the-art (SOTA) performance on five mainstream video segmentation benchmarks.
@@ 1 Introduction
However, these query-based methods exhibit a notable limitation: they tend to underperform when dealing with newly emerging or disappearing objects, even with recent state-of-the-art (SOTA) methods [73, 21]. As shown in the left part of Fig. 1, these methods only achieve less than 45% recall ratio on emerging and disappearing objects. As illustrated in the middle part of Fig. 1, they often fail to detect newly emerging objects. And they may incorrectly track a different, similar object when the original object disappears.
@@ 2 Related Work
Video Instance Segmentation (VIS). Earlier works [68, 35, 69] extend image instance segmentation methods by adding tracking heads and learning the feature association. With the rise of vision transformer [8, 55, 16], current SOTA video instance methods [55, 22, 10, 60, 23, 61, 70, 33, 37, 28, 18] adopt query-based designs. In particular, Video K-Net [34] and IDOL [61] directly learn the query association embeddings via contrastive learning. After that, association-based methods [68, 23, 61, 70, 33, 37], accomplish the tracking of video targets between adjacent frames or adjacent video clips through object matching. Meanwhile, several models [22, 10, 60] directly predict 3D volumes of video instances across times in a semi-online manner. On the other hand, propagation-based methods [21, 28, 18, 73] delegate the tracking problem to the transformer decoder. They take the object queries outputted from the previous frame or video clip as input and iteratively optimize them to segment the objects in the current frame. Inspired by the design of the dual encoder in the low-level area [53] and multi-task learning [31, 63], several works have also decoupled the tracking decoder and segmentation decoder. Note that these methods can also be applied for video panoptic segmentation (VPS) and video semantic segmentation (VSS) tasks [48, 40, 39, 56, 24, 80] by adding queries for segmenting the stuff. However, both poorly handle challenges such as new-emerging objects and target disappearance. Our approach improves the model’s ability to address these challenges by employing dynamic anchor queries to manage the emergence and disappearance of objects.
@@
Universal Video Segmentation. Current efforts in the field of video segmentation directly draw upon the design of universal image segmentation models [57, 75, 10, 12, 66, 13, 54] to develop universal video segmentation models [34, 26, 33, 1, 73, 30, 59], achieving comparable or even superior performance to specialized video segmentation methods. Video K-Net [34] and TubeFormer [26] are the first to unify all three video segmentation tasks. TarVIS [1] extends the concept of mask classification from image segmentation to video segmentation, accomplishing universal video segmentation via one shared model. DVIS [73] receives frame object query outputs from Mask2Former [10] for universal video segmentation. Recently, several works [32, 64] unified both image and video segmentation in one model. However, these methods are still unaware of object emergence and disappearance problems, which are significant for long videos.
@@
Video Object Tracking. Object tracking [72, 50, 43, 76, 51, 65, 42, 17, 14, 7, 3, 77] is a crucial task in VIS and VPS [78, 24, 25, 4, 29, 19], and many works adopt the tracking-by-detection paradigm [6, 27, 62, 79, 46, 58, 52, 38]. These methods divide the task into two sub-tasks, where an object detector first detects objects and then associates them using a tracking algorithm. Recent works [44, 15, 49, 2] also perform clip-wise segmentation and tracking. However, the former [44, 15] only focuses on single-object mask tracking, while the latter [49] considers global tracking via clip-level matching.
@@ 4.3 Ablation Studies and Analysis
Qualitative results comparison. In Fig. 7, we present the visual comparison with our baseline. In this challenging scene, there are new-emerging, disappearing, and reappearing video objects. Our method accurately addresses all these scenarios, whereas DVIS suffers from issues including missed new-emerging objects and ID switches.
@@ 0.A.4 More Qualitative Results
Fig. A1 presents qualitative results in complex scenes. To fully demonstrate the performance of our method in challenging scenes such as fast motion and occlusions, we also provide supplementary videos for reference at project page.